OpenStudy (anonymous):

If a number is increased by its square, the result is 72. Find the number.

OpenStudy (kainui):

Consider writing out exactly that in an equation form as "a number" being "x". "the result is" as "="

OpenStudy (anonymous):

let x be that number a number increased by its square is x + x^2 = 72 x^2 +x - 72 = 0 (x+9)(x-8) = 0 equate each factor to zero then solve for x

OpenStudy (anonymous):

x+x^2=72 sfx

OpenStudy (shane_b):

@mathwhiz99 has already explained it pretty well.

OpenStudy (anonymous):

Yup, he did!

OpenStudy (anonymous):

If a number is increased by its square, the result is 72. Find the number. ( x ) ( + ) ( x^2 ) ( = 72) x+x^2=72

OpenStudy (anonymous):

the answer is 8

OpenStudy (shane_b):

Congratulations.

OpenStudy (anonymous):

and -9